OpenStudy (anonymous):

Verify the identity. sin(3pi - x) = sinx

OpenStudy (anonymous):

Remember that Sin(x+y) = Sin(x)Cos(y) - Sin(y)Cos(x) Knowing that, can't we just call the whole thing: Sin(3pi + (-x)) From there, we apply the identity Sin(3pi)Cos(x) - Sin(x)Cos(3pi) We know that Sin(3pi) = 0, and Cos(3pi) = -1 Sin(3pi)Cos(x) - Sin(x)Cos(3pi) = (0) Cos(x) - Sin(x) (-1) Anything times zero is zero, and two negatives make a positive so: = 0 + Sin(x) = Sin(x)

OpenStudy (anonymous):

oh hang on it should be -x and it should be Sin(3pi)Cos(x) + Sin(x)Cos(3pi)

OpenStudy (anonymous):

yeah i understand. all i was confused about was how to simplify further after using the sum and difference formula things. Thanks!